Find values for m and b in the following system so that the solution to the system
is (1, 2).
Mx + 3y= 8
-3x+ 4y+ b
Thank you.

Answer by malakumar_kos@yahoo.com(315)


Find values for m and b in the following system so that the solution to the system
is (1, 2).
Mx + 3y= 8
-3x+ 4y+ b
since the solutions of the system are given it means values of x&y are known by substituting for x&y we can find the values of m ^b
substitute for x &y in Mx+3y = 8 we get M(1)+3(2) = 8 or M+6 = 8

M = 8-6 = 2
-3(1)+4(2) = b or -3+8 = b hence b = 5
the values of M = 2 and b = 5.